A rock group plans three concert tours over a period of 38 weeks. The tour in Britain will be 4 weeks longer than the tour in France and the tour in Germany will be 2 weeks shorter than the tour in France. How many weeks will they be in France?
Answer by stanbon(75887)

A rock group plans three concert tours over a period of 38 weeks. The tour in Britain will be 4 weeks longer than the tour in France and the tour in Germany will be 2 weeks shorter than the tour in France. How many weeks will they be in France?
---------------
Equations:
b = f + 4
g = f-2
b + f + g = 18
-----
Substitute for "b" and for "g" and solve for "f":
f+4 + f + f-2 = 18
3f + 2 = 18
3f = 16
f = 5 1/3 weeks (France)
b = f+4 = 9 1/3 weeks (Britain)
g = f-2 = 3 1/3 weeks (Germany)
===============
Cheers,
Stan H.
==================